Circular No. 7079

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA MUSCAE 1998 W. Liller, Viña del Mar, Chile, forwards the following position for the

star reported on IAUC 7078, measured from a CCD frame taken with a 0.20-m f/1.5 Schmidt camera: α = 11^{h}56^{m}08^{s}.9, δ = −65^{o}34′20′′.6 (equinox 2000.0; standard deviations 0^{s}.2 and 0′′.4). An objective-prism spectrum taken with the same instrument shows Hα to be exceptionally bright (∼ 7.3 times brighter than the neighboring continuum); Hβ is weakly present. On Dec. 1.1754 UT, V = 9.55.

B. M. Ewen-Smith, Nuriootpa, South Australia, reports the following visual magnitude estimate: Dec. 31.495 UT, 9.9.

COMET C/1998 U5 (LINEAR)

Ephemeris extension to IAUC 7059 (elements MPC 33189):

1998 /99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 13 21^{h}26.^{m}73 +25^{o}22.^{′}1 1.013 1.244 77.^{o}0 50.^{o}5  9.0
23 21 16.09 +20 27.3 1.292 1.236  64.1 45.7  9.5
Jan. 2 21 11.11 +17 33.8 1.551 1.248  53.5 39.3  9.9
12 21 08.70 +15 45.6 1.781 1.278  44.3 32.5 10.3
22 21 07.60 +14 37.5 1.974 1.324  36.8 26.4 10.7
Feb. 1 21 07.10 +13 56.8 2.129 1.385  31.3 21.7 11.1
11 21 06.74 +13 36.3 2.243 1.458  28.8 19.0 11.4
21 21 06.15 +13 31.9 2.318 1.541  29.8 18.6 11.7
Mar. 3 21 04.93 +13 40.7 2.355 1.630  34.0 19.9 12.0
13 21 02.66 +14 00.3 2.357 1.726  40.5 22.0 12.2
23 20 58.89 +14 28.9 2.327 1.826  48.5 24.1 12.4
Apr. 2 20 53.04 +15 04.1 2.271 1.928  57.5 25.9 12.6
12 20 44.48 +15 42.9 2.194 2.033  67.5 27.1 12.8
22 20 32.47 +16 21.0 2.103 2.140  78.3 27.4 12.9
May 2 20 16.24 +16 51.8 2.009 2.247  90.0 26.6 13.0
12 19 55.24 +17 06.0 1.922 2.355 102.4 24.8 13.1

Total visual magnitude and coma-diameter estimates: Nov. 26.42 UT, 8.4, 10′ (N. Biver, Hale-Pohaku, HI, 7×50 binoculars); Dec. 6.84, 9.1, 5′ (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars); 14.12, 8.8, 8′ (R. Keen, Mt. Thorodin, CO, 0.15-m reflector); 18.72, 9.6, 6′ (M. Meyer, Frauenstein, Germany, 20×100 binoculars); 24.73, 9.1, 3′ (A. Diepvens, Balen, Belgium, 0.15-m refractor); 30.70, 10.8, 4′.0 (M. Lehký, Hradec Králové, Czech Republic, 25×100 binoculars).

1998 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green